Are the following lines parallel, perpendicular, or neither? L1 through (9, 2) and (3, –8) L2 through (–3, 5) and (5, –1)
Answer by stanbon(75887)

Are the following lines parallel, perpendicular, or neither? L1 through (9, 2) and (3, –8) L2 through (–3, 5) and (5, –1)
------------
Slope of L1 = [2--8]/[9-3]=10/6=5/3
Slope of L2 = [-1-5]/[5--3]= -6/8=-3/4
-------------
L1 and L2 are not parallel or perpendicular.
Cheers,
Stan H.